# Kernel Banzhaf: A Fast and Robust Estimator for Banzhaf Values

### v1

@@ 3.3 Approximation Guarantees
Theorem 3.3 is a standard guarantee for leverage score sampling. However, establishing the theorem for Kernel Banzhaf requires completely reproving it from scratch because of the incorporation of paired sampling in Kernel Banzhaf. For the analysis of Leverage SHAP, Musco & Witter (2024) prove a similar theorem with paired leverage score sampling but using different techniques, since the samples in their algorithm are taken without replacement.
The theorem can only be improved in the logarithmic factor and dependence on $\delta$ and $\epsilon$. To see why, consider a linear set function where $\mathbf{b}$ is in the span of $\mathbf{A}$. Then the right-hand side is 0 and we must recover $\bm{\phi}$ exactly. To do this, we need to observe at least $n$ linearly independent rows hence we need at least $\Omega(n)$ samples.
While it gives a strong guarantee, the term in Theorem 3.3 is less interpretable. Fortunately, we can use the special properties of $\mathbf{A}$ and $\mathbf{b}$ to prove the following corollary in terms of the $\ell_{2}$-norm error.
Corollary 3.4.
Define $\gamma=\|\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}/\|\mathbf{A}\bm{\phi}\|_{2}^{2}$. Any $\hat{\bm{\phi}}$ that satisfies Equation 10 also satisfies
$\|\hat{\bm{\phi}}-\bm{\phi}\|_{2}^{2}\leq\epsilon\gamma\|\bm{\phi}\|_{2}^{2}.$ (11)
The proof is similar to guarantees for general linear regression problems and for the Leverage SHAP setting (see e.g., Lemma 68 in Drineas & Mahoney (2018) and Corollary 4.1 in Musco & Witter (2024)). However, because of the special structure of our problem, our result differs in that there is no relaxation in the guarantee from Theorem 3.3 to Corollary 3.4. Hence the (near-)optimality of Theorem 3.3 implies the (near-)optimality of Corollary 3.4.
Notice that Theorem 3.3 has no dependence on $\gamma$ while Corollary 3.4 does. This suggests that, when $\mathbf{b}$ is outside the span of $\mathbf{A}$, Kernel Banzhaf can recover a solution $\hat{\bm{\phi}}$ that has near optimal objective value but is far from the optimal solution $\bm{\phi}$. The reason is that there can be many (equally bad) solutions for the optimal vector that recover similar objective values but are far from each other. The benefit of the dependence on $\gamma$ is that we recover an even stronger guarantee when $\mathbf{b}$ is in the span of $\mathbf{A}$. In fact, we recover the optimal solution exactly since $\gamma=0$ in this case.
@@
$\|\mathbf{A}\hat{\bm{\phi}}-\mathbf{b}\|_{2}^{2}=\|\mathbf{A}\hat{\bm{\phi}}-\mathbf{A}\bm{\phi}+\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}=\|\mathbf{A}\hat{\bm{\phi}}-\mathbf{A}\bm{\phi}\|_{2}^{2}+\|\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}$ (12)
where the second equality follows because $\mathbf{A}\bm{\phi}-\mathbf{b}$ is orthogonal to any vector in the span of $\mathbf{A}$. To see why this orthogonality holds, consider that $\bm{\phi}$ minimizes $\|\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}$, implying $\mathbf{A}^{T}(\mathbf{A}\bm{\phi}-\mathbf{b})=\mathbf{0}$. From the zero gradient condition,
$(\mathbf{A}\hat{\bm{\phi}}-\mathbf{A}\bm{\phi})^{T}(\mathbf{A}\bm{\phi}-\mathbf{b})=(\hat{\bm{\phi}}-\bm{\phi})^{T}\mathbf{A}^{T}(\mathbf{A}\bm{\phi}-\mathbf{b})=(\hat{\bm{\phi}}-\bm{\phi})^{T}\mathbf{0}=0,$ (13)
which confirms that $\mathbf{A}\bm{\phi}-\mathbf{b}$ is orthogonal to $\mathbf{A}\hat{\bm{\phi}}-\mathbf{A}\bm{\phi}$.
Then by assumption, we have $\|\mathbf{A}\hat{\bm{\phi}}-\mathbf{A}\bm{\phi}\|_{2}^{2}\leq\epsilon\|\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}$. By Observation 3.1, we have
$\|\mathbf{A}\bm{\phi}\|_{2}^{2}=\bm{\phi}^{\top}\mathbf{A^{\top}A}\bm{\phi}=2^{n-2}\|\bm{\phi}\|_{2}^{2}$ (14)
and, similarly, $\|\mathbf{A}(\hat{\bm{\phi}}-\bm{\phi})\ \|_{2}^{2}=2^{n-2}\|\hat{\bm{\phi}}-\bm{\phi}\ \|_{2}^{2}$. Then, with the definition of $\gamma$,
$2^{n-2}\|\hat{\bm{\phi}}-\bm{\phi}\|_{2}^{2}=\|\mathbf{A}(\hat{\bm{\phi}}-\bm{\phi})\|_{2}^{2}\leq\epsilon\|\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}=\epsilon\gamma\|\mathbf{A}\bm{\phi}\|_{2}^{2}=2^{n-2}\epsilon\gamma\|\bm{\phi}\|_{2}^{2}.$ (15)
@@
Because of the structure of $\mathbf{A}$, Theorem 3.3 and Corollary 3.4 are equivalent. Therefore, since we believe that Theorem 3.3 is nearly optimal, we believe that $\gamma$ is a necessary term in the error bound of Corollary 3.4.
Additionally, we emphasize that Kernel Banzhaf and its theoretical guarantees apply to Banzhaf values defined on any set function.
Figure 2: Plots comparing the $\ell_{2}$-norm errors (i.e. $\|\hat{\bm{\phi}}-\bm{\phi}\|_{2}^{2}$) for Kernel Banzhaf (and its ablated version without paired sampling), MC, and MSR across increasing sample sizes in eight datasets. Each point represents the median of 50 runs, with shaded areas indicating the 25th to 75th percentiles. The plots showcase the effectiveness of Kernel Banzhaf in various real-world datasets.
4 Experiments
We conduct a detailed experimental evaluation and compare Kernel Banzhaf against state-of-the-art estimators across eight popular datasets with varying size (having between 8 and 241 features)  [^3] Our code is available at https://github.com/lyrain2001/kernel-banzhaf. These datasets have been used in prior research for semi-value-based model explanation (Lundberg & Lee, 2017; Covert & Lee, 2020; Lundberg et al., 2020; Karczmarz et al., 2022), and they are described in detail in Appendix C. To assess the accuracy of the estimators, we employ the $\ell_{2}$-norm error as our primary error metric for evaluation. While prior work used convergence ratio as a measure of effectiveness (Wang & Jia, 2023; Covert & Lee, 2020), convergence is insufficient to measure how close the estimators are to exact Banzhaf values. We use XGBoost (Chen & Guestrin, 2016) for our models in the main experiments, which makes it possible to apply the tree-based algorithm of (Karczmarz et al., 2022) to compute exact Banzhaf values. Implementation details are given in Appendix B. We also assess the effectiveness of our approach on neural network models using smaller datasets for which Equation 2 can be used to calculate exact Banzhaf values. These experiments are presented in Appendix D and their results are consistent with those for the tree models. In addition to $\ell_{2}$-norm, we use the objective value naturally suggested by the linear regression formulation to measure the estimation error; the results in Appendix E corroborate the superior performance of Kernel Banzhaf.
4.1 Comparing Banzhaf Estimators
We compare Kernel Banzhaf with the following methods for estimating Banzhaf values:
MC The Monte Carlo (MC) algorithm estimates each Banzhaf value individually. Let $\mathcal{S}_{i}$ be the subsets sampled for player $i\in[n]$. We have that $\sum_{i}|\mathcal{S}_{i}|=m$. The MC estimate for player $i$ is given by
$\hat{\phi}^{\textrm{MC}}_{i}=\frac{1}{|\mathcal{S}_{i}|}\sum_{S\in\mathcal{S}_{i}}[v(S\cup\{i\})-v(S)].$
The disadvantage of this algorithm is that each sample is used only for a single player.
$\hat{\phi}^{\textrm{MSR}}_{i}=\frac{1}{|\mathcal{S}_{\ni i}|}\sum_{S\in\mathcal{S}_{\ni i}}v(S)-\frac{1}{|\mathcal{S}_{\not\owns i}|}\sum_{S\in\mathcal{S}_{\not\owns i}}v(S).$
Kernel Banzhaf (Excluding Pairs) This algorithm is identical to Kernel Banzhaf (Algorithm 1) but without paired sampling.
As we subsample feature sets for estimating Banzhaf value, sample size is an important concern in real-world settings. Our first experiment investigates the error of each estimator by number of samples. As shown in Figure 2, Kernel Banzhaf outperforms the other algorithms for all datasets over different sample sizes. As expected, the error for all estimators decreases as the sample size increases. Interestingly, MSR performs worse than MC when applied to datasets with a large number of features (e.g., the NHANES dataset and larger), particularly as the number of samples increases. We believe this is because the MSR algorithm has high variance which is related to the generally large magnitude of the set function. In prior work, the MSR algorithm was shown to converge faster than the MC algorithm, but faster convergence does not always imply better accuracy.
The set functions in explainable AI tasks, being based on stochastically-trained models, often exhibit variance. We investigated how the estimation errors vary with random noise added to the set function, and as shown in Figure 3, Kernel Banzhaf and its variant without paired sampling are more robust to this noise than the other algorithms, especially when the level of noise size is low. This suggests that Kernel Banzhaf is particularly well-suited for real-world applications where the set function is approximated.
Figure 3: Plots of $\ell_{2}$-norm error by noise levels across Banzhaf estimators. For each noise level $\sigma$, the estimator observes $v(S)+x$ where $x\sim N(0,\sigma)$. Kernel Banzhaf outperforms for lower noise levels, eventually matching its ablated version and MSR for larger noise. MC is worse for all noise settings, likely because its constituent estimates are $v(S\cup\{i\})+x-v(S)-x^{\prime}$, essentially doubling the noise variance.
4.2 Comparing Kernel and Shapley Regressions
Prior work found that Banzhaf values can be estimated more efficiently and robustly than Shapley values (Karczmarz et al., 2022; Wang & Jia, 2023). To this end, we compare Kernel Banzhaf with state-of-the-art algorithms for estimating Shapley values: Optimized KernelSHAP, which employs paired sampling and sampling without replacement techniques (Lundberg & Lee, 2017; Covert & Lee, 2020) and Leverage SHAP, which employs leverage score sampling in addition to the aforementioned techniques (Musco & Witter, 2024). To obtain ground-truth Shapley values, we use the interventional setting of Tree SHAP (Lundberg et al., 2020). We employ the normalized $\ell_{2}$-norm squared error, defined as $\|\hat{\bm{\phi}}-\bm{\phi}\|_{2}^{2}/\|\bm{\phi}\|_{2}^{2}$ for comparison to ensure fairness.
When we compare the estimation error by varying the number of samples (Figure 9, Appendix F), we observe that for larger $n$, Kernel Banzhaf gives the best performance, followed by Leverage SHAP then Optimized KernelSHAP. Note that the larger the number of features, the bigger the difference in estimation error. When $n$ is small as in the Diabetes dataset, KernelSHAP and Leverage SHAP perform better because they sample without replacement.
We also investigate the impact of noise. To ensure parity between the Banzhaf and Shapley value estimators, we normalize the noise based on the raw output of the set function. The results are shown in Figure 4. The initial data points represent scenarios with no added noise, so our focus is primarily on the trends of the lines. The horizontal line representing Kernel Banzhaf, which remains unchanged as noise levels increase, underscores the robustness of Kernel Banzhaf compared to the Shapley value estimators.
Since Kernel Banzhaf is similar in design to KernelSHAP and Leverage SHAP, at first sight, it is surprising that Kernel Banzhaf performs better for large $n$. For both Kernel Banzhaf and Leverage SHAP, a crucial component of the analysis is to build the subsampled matrix $\tilde{\mathbf{A}}$ so that
$c\mathbf{A}^{\top}\mathbf{A}\preceq\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}\preceq C\mathbf{A}^{\top}\mathbf{A}$ (16)
where $c\leq C$ are two close scalars and $\preceq$ denotes the Loewner order. Multiplying on the left and right by $(\mathbf{A}^{\top}\mathbf{A})^{-\frac{1}{2}}$ yields
$c\mathbf{I}\preceq(\mathbf{A}^{\top}\mathbf{A})^{-\frac{1}{2}}\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}(\mathbf{A}^{\top}\mathbf{A})^{-\frac{1}{2}}\preceq C\mathbf{I}.$ (17)
Let $\mathbf{K}=(\mathbf{A}^{\top}\mathbf{A})^{-\frac{1}{2}}\tilde{\mathbf{A}}^{\top}\tilde{\mathbf{A}}(\mathbf{A}^{\top}\mathbf{A})^{\frac{1}{2}}$. Therefore, we can measure how close the best $C$ and $c$ are by computing the condition number of $\mathbf{K}$; a lower condition number implies $C$ is closer to $c$ and hence the subsampling preserves the important properties of the regression problem. Figure 5 plots the condition number of this matrix for all three regression-based methods; the plots suggest that Kernel Banzhaf constructs a better subproblem at all scales, explaining its superior performance.
Figure 4: Normalized $\ell_{2}$-norm error (i.e. $\|\hat{\bm{\phi}}-\bm{\phi}\|_{2}^{2}/\|\bm{\phi}\|_{2}^{2}$) by noise level for Banzhaf and KernelSHAP estimators. This plot displays the $\ell_{2}$-norm error as a function of noise level across eight datasets similar to Figure 3. Note that a more horizontal trend line is preferable. Kernel Banzhaf consistently exhibits superior performance by maintaining lower error rates amid increasing noise, compared to its competitors whose error rates rise. This demonstrates Kernel Banzhaf’s robustness and efficiency, marked by its minimal loss of precision.
5 Related Work
Game-Theoretic Explainable AI
Semivalues like Shapley and Banzhaf values, derived from game theory, are widely used for machine learning tasks. Feature attribution involves determining the impact of individual features on model predictions, with Shapley values notably used in Lundberg & Lee (2017). This method also facilitates feature selection by modifying set functions to measure the feature’s impact on model performance via loss metrics (Covert et al., 2020). Data valuation extends to treat individual data points as players, assessing their influence on model training outcomes like accuracy or AUC (Ghorbani & Zou, 2019).
A parallel trend in utilizing Banzhaf values has been noted, with studies indicating their robustness compared to Shapley values, particularly under conditions of low numerical precision (Karczmarz et al., 2022). Furthermore, Wang & Jia (2023) established the superiority of Banzhaf values among semivalues regarding safety margin, the threshold of set function perturbations that the value order can withstand without changing. These properties have led to extensive application of Banzhaf values for feature attribution (Datta et al., 2015; Kulynych & Troncoso, 2017; Sliwinski et al., 2018; Patel et al., 2021; Karczmarz et al., 2022) and, more recently, in data valuation, where Wang & Jia (2023) and Li & Yu (2024) have shown their superior performance compared to Shapley values.
Shapley Estimators
KernelSHAP, introduced by Lundberg & Lee (2017), employs game-theoretic Shapley values in a linear regression framework to explain any machine learning model’s predictions. The method’s efficiency and applicability are enhanced through paired sampling and sampling without replacement, which balances the number of samples used to update each Shapley value, as refined by Covert & Lee (2020); Jethani et al. (2021). In recent work, Musco & Witter (2024) proposed Leverage SHAP, a variant of KernelSHAP that uses leverage score sampling to build the subsampled linear regression problem. Leverage scores are a statistical quantity that intuitively measure the importance of a data point (Woodruff et al., 2014; Drineas & Mahoney, 2018). Leverage SHAP outperforms KernelSHAP and, because of the mathematical properties of leverage scores, provably returns accurate approximations. Meanwhile, TreeSHAP, developed by Lundberg et al. (2020), specifically tailors the SHAP framework for tree-based models, ensuring faster, exact computations by directly incorporating tree structures into Shapley value calculations.
Banzhaf Estimators
Merrill III (1982) initially introduced approximating the Banzhaf power index through Monte Carlo sampling across subsets. Building on this, Bachrach et al. (2010) applied Hoeffding bounds to analyze the sample complexity required for these approximations. Further advancing the method, Wang & Jia (2023) developed the Maximum Sample Reuse (MSR) Monte Carlo estimator. This approach samples sets individually for each term in the marginal utility calculation, facilitating the reuse of samples across different computations of marginal utility. However, due to this variation, the MSR estimates of Banzhaf values are correlated among different players and exhibit large variance. Moreover, for structured set functions such as decision trees, where predictions are aggregated from root to leaf, Banzhaf values can be computed with exact precision (Karczmarz et al., 2022). This computational efficiency is attributable to the ability to independently isolate and assess the impact of each feature along the decision path. Despite its effectiveness within this context, this method lacks generalizability to other models and machine learning tasks.
Banzhaf Values and Regression
Hammer & Holzman (1992) shows how least squares regression can approximate the Banzhaf power index, which only requires a ‘simple’ set function $v:2^{N}\rightarrow\{0,1\}$. Their regression formulation differs from what we propose in Theorem 3.2 and may not yield exact Banzhaf values even when no subsampling is done. Alonso-Meijide et al. (2015) proposed modifications of Banzhaf values to enable them to satisfy the Efficiency property, via additive and multiplicative normalization, and formulated a least squares regression to obtain them.
Figure 5: Plot of condition number of matrix $\mathbf{K}$ as defined in Equation 17 by sample size for estimators based on linear regression in the context of Banzhaf and Shapley value estimations. This plot demonstrates the numerical stability of the Kernel Banzhaf estimator by showing its lower condition numbers across varying sample sizes, highlighting its superior robustness and reduced sensitivity to numerical errors as shown in Figure 4, particularly when $n$ is large.
6 Conclusion
In this paper, we presented Kernel Banzhaf, an innovative estimator for Banzhaf values that leverages linear regression to improve the efficiency and robustness of Banzhaf value estimation for explainable AI. Our empirical and theoretical analysis reveal that Kernel Banzhaf outperforms existing methods for Banzhaf estimation. Additionally, our findings highlight its superior robustness compared to Shapley estimators, establishing Kernel Banzhaf value as a viable and effective alternative to other semivalue-based methods for feature explanation in machine learning models.
Future research can focus on extending Kernel Banzhaf to diverse set functions and examine its potential for integration into real-time systems. This will help us to further understand the scalability and applicability of our approach in diverse machine learning contexts. Additionally, a concrete direction for future work is to develop adaptations of Kernel Banzhaf for weighted Banzhaf values (Li & Yu, 2024).
Acknowledgements
Freire and Liu were partially supported by NSF awards IIS-2106888 and the DARPA ASKEM and ARPA-H BDF programs. Liu initiated this project during her summer internship at Google Research. Witter was supported by the National Science Foundation under Graduate Research Fellowship Grant No. DGE-2234660. Opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not reflect the views of NSF or DARPA. We thank Christopher Musco for his invaluable contribution to the problem formulation and experiment design.
Appendix A Theoretical Guarantees
See 3.3
@@ Appendix A Theoretical Guarantees
Theorem 3.3 is a standard guarantee for leverage score sampling. However, because rows are sampled in pairs, we need to substantially modify the standard analysis. In particular, both the spectral guarantee that the sampling matrix preserves eigenvalues and the Frobenius guarantee that the sampling matrix preserves Frobenius norm need to be reproved.
@@
We will adopt the notation from Wu (2018); Musco & Witter (2024). Let’s consider a leverage score sampling method where rows are selected in blocks. Define $\Theta$ as a partition of blocks, each containing 2 elements with identical leverage scores in our setting. We assign a sampling probability $p^{+}_{i}$ to each block $\Theta_{i}$, calculated as the sum of leverage scores in that block divided by the total sum of all leverage scores: $p^{+}i\vcentcolon=\frac{\sum{k\in\Theta_{i}}\ell_{k}}{\sum_{j}\ell_{j}}$. For simplicity of notation, suppose $m$ is even. Let $\mathbf{S}\in\mathbb{R}^{m\times\rho}$ represent our random sampling matrix, initially filled with zeros. The sampling process repeats $m/2$ times: Sample a block $\Theta_{i}$ with probability $p^{+}_{i}$. For each $k\in\Theta_{i}$, set the $k$th entry in an empty row to $\frac{1}{\sqrt{p^{+}_{i}}}$.
@@
Fact A.2 (Matrix Chernoff).
@@
$\mathbb{E}[\mathbf{X}_{j}]=\mathbf{I}-\sum_{i=1}^{|\Theta|}p^{+}_{i}\frac{1}{p^{+}_{i}}\sum_{k\in\Theta_{i}}\mathbf{U}_{k}^{\top}\mathbf{U}_{k}=\mathbf{0}$ (21)
@@
$\mathbb{E}[\mathbf{X}_{j}^{2}]$ $=\mathbf{I}-2\sum_{i=1}^{|\Theta|}p^{+}_{i}\frac{\sum_{k\in\Theta_{i}}\mathbf{U}_{k}^{\top}\mathbf{U}_{k}}{p^{+}_{i}}+\sum_{i=1}^{|\Theta|}p^{+}_{i}\frac{\left(\mathbf{U}_{\Theta_{i}}^{\top}\mathbf{U}_{\Theta_{i}}\right)^{2}}{{p^{+}_{i}}^{2}}$ (23)
@@
(24$)\preceq n\sum_{i=1}^{|\Theta_{i}|}\frac{2\ell^{\max}_{i}|\Theta_{i}|^{2}\mathbf{U}_{\Theta_{i}}^{\top}\mathbf{U}_{\Theta_{i}}}{\sum_{k\in\Theta_{i}}\ell_{k}}\preceq n\max_{i}2|\Theta_{i}|\frac{\ell^{\max}_{i}}{\ell^{\min}_{i}}\mathbf{I}.$ (25)
Since $|\Theta_{i}|\leq 2$ and the leverage scores in a block are all equal, $\|\mathbb{E}[\mathbf{X}_{j}^{2}]\|_{2}\leq 4n$.
Applying Fact A.2 with $m=O(n\log(n/\delta)/\epsilon^{2})$ yields
@@
$\Pr\left(\left\|\mathbf{U^{\top}S^{\top}SV}-\mathbf{U^{\top}V}\right\|_{F}>\epsilon\|\mathbf{U}\|_{F}\|\mathbf{V}\|_{F}\right)$ $\leq\frac{\mathbb{E}\left[\left\|\mathbf{U^{\top}S^{\top}SV}-\mathbf{U^{\top}V}\right\|_{F}^{2}\right]}{\epsilon^{2}\|\mathbf{U}\|_{F}^{2}\|\mathbf{V}\|_{F}^{2}}\leq\frac{1}{m\epsilon^{2}}\leq\delta$ (30)
@@
With Lemmas A.1 and A.3 already proved for the special paired leverage score sampling, the following analysis is standard. For example, see Appendix A in Musco & Witter (2024) which proves the same guarantee but for sampling without replacement, requiring yet different proofs of the spectral and Frobenius approximation lemmas. We include the following proof for completeness.
Proof of Theorem 3.3.
@@
$\|\mathbf{A}\tilde{\bm{\phi}}-\mathbf{b}\|_{2}^{2}=\|\mathbf{A}\tilde{\bm{\phi}}-\mathbf{A}\bm{\phi}+\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}=\|\mathbf{A}\tilde{\bm{\phi}}-\mathbf{A}\bm{\phi}\|_{2}^{2}+\|\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}$ (31)
where the second equality follows because $\mathbf{A}\bm{\phi}-\mathbf{b}$ is orthogonal to any vector in the span of $\mathbf{A}$. So to prove the theorem, it suffices to show that
$\|\mathbf{A\tilde{\bm{\phi}}-A\bm{\phi}}\|_{2}^{2}\leq\epsilon\|\mathbf{A\bm{\phi}-b}\|_{2}^{2}.$ (32)
Let $\mathbf{U}\in\mathbb{R}^{\rho\times n}$ be an orthonormal matrix that spans the columns of $\mathbf{A}$. There is some $\mathbf{y}$ such that $\mathbf{Uy}=\mathbf{A}\bm{\phi}$ and some $\tilde{\mathbf{y}}$ such that $\mathbf{U}\tilde{\mathbf{y}}=\mathbf{A}\tilde{\bm{\phi}}$. Observe that $\|\mathbf{A\tilde{\bm{\phi}}-A\bm{\phi}}\|_{2}=\|\mathbf{U}\tilde{\mathbf{y}}-\mathbf{Uy}\|_{2}=\|\tilde{\mathbf{y}}-\mathbf{y}\|_{2}$ where the last equality follows because $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}$.

### v2

@@ Appendix A Kernel Banzhaf Approximation Guarantees
In order to prove Theorem 3.2, we will show that uniform sampling is actually equivalent to leverage score sampling. The leverage scores of a matrix are an important statistical quantity that roughly measures the “uniqueness” of a row (Woodruff et al., 2014; Drineas & Mahoney, 2018).
The leverage scores of $\mathbf{A}$ are given by
$\ell_{S}=[\mathbf{A}]_{S}^{\top}\left(\mathbf{A^{\top}A}\right)^{-1}[\mathbf{A}]_{S}=\frac{1}{2^{n-2}}\left(\mathbbm{1}[S]-\frac{1}{2}\right)^{\top}\left(\mathbbm{1}[S]-\frac{1}{2}\right)=\frac{n}{2^{n}}$ (12)
where the second equality follows by Equation 6. Notice the leverage scores are equivalent to uniform sampling and Theorem 3.2 immediately follows.
We can now prove Corollary 3.3.
See 3.3
@@
$\|\mathbf{A}\tilde{\boldsymbol{\phi}}-\mathbf{b}\|_{2}^{2}=\|\mathbf{A}\tilde{\boldsymbol{\phi}}-\mathbf{A}\boldsymbol{\phi}+\mathbf{A}\boldsymbol{\phi}-\mathbf{b}\|_{2}^{2}=\|\mathbf{A}\tilde{\boldsymbol{\phi}}-\mathbf{A}\boldsymbol{\phi}\|_{2}^{2}+\|\mathbf{A}\boldsymbol{\phi}-\mathbf{b}\|_{2}^{2}$ (13)
where the second equality follows because $\mathbf{A}\boldsymbol{\phi}-\mathbf{b}$ is orthogonal to any vector in the span of $\mathbf{A}$. Then, by the assumption that $\|\mathbf{A}\tilde{\boldsymbol{\phi}}-\mathbf{b}\|_{2}^{2}\leq(1+\epsilon)\|\mathbf{A}\boldsymbol{\phi}-\mathbf{b}\|_{2}^{2}$, we have $\|\mathbf{A}\tilde{\boldsymbol{\phi}}-\mathbf{A}\boldsymbol{\phi}\|_{2}^{2}\leq\epsilon\|\mathbf{A}\boldsymbol{\phi}-\mathbf{b}\|_{2}^{2}$. By Equation 6, we have
$\|\mathbf{A}\boldsymbol{\phi}\|_{2}^{2}=\boldsymbol{\phi}^{\top}\mathbf{A^{\top}A}\boldsymbol{\phi}=2^{n-2}\|\boldsymbol{\phi}\|_{2}^{2}$ (14)
and, similarly, $\|\mathbf{A}(\tilde{\boldsymbol{\phi}}-\boldsymbol{\phi})\ \|_{2}^{2}=2^{n-2}\|\tilde{\boldsymbol{\phi}}-\boldsymbol{\phi}\ \|_{2}^{2}$. Then, with the definition of $\gamma$,
$2^{n-2}\|\tilde{\boldsymbol{\phi}}-\boldsymbol{\phi}\|_{2}^{2}=\|\mathbf{A}(\tilde{\boldsymbol{\phi}}-\boldsymbol{\phi})\|_{2}^{2}\leq\epsilon\|\mathbf{A}\boldsymbol{\phi}-\mathbf{b}\|_{2}^{2}=\epsilon\gamma\|\mathbf{A}\boldsymbol{\phi}\|_{2}^{2}=2^{n-2}\epsilon\gamma\|\boldsymbol{\phi}\|_{2}^{2}.$ (15)
@@
Next, we prove Theorem A.1, which is a standard guarantee for leverage score sampling. However, because rows are sampled in pairs, we need to modify the analysis. In particular, both the spectral guarantee that the sampling matrix preserves eigenvalues and the Frobenius guarantee that the sampling matrix preserves Frobenius norm need to be reproved.
@@
We will adopt the notation from Wu (2018). Let’s consider a leverage score sampling method where rows are selected in blocks. Define $\Theta$ as a partition of blocks, each containing 2 elements with identical leverage scores in our setting. We assign a sampling probability $p^{+}_{i}$ to each block $\Theta_{i}$, calculated as the sum of leverage scores in that block divided by the total sum of all leverage scores: $p^{+}i\vcentcolon=\frac{\sum{k\in\Theta_{i}}\ell_{k}}{\sum_{j}\ell_{j}}$. For simplicity of notation, suppose $m$ is even. Let $\mathbf{S}\in\mathbb{R}^{m\times\rho}$ be a sampling matrix, initially set to $\mathbf{0}_{m\times n}$. The sampling process repeats $m/2$ times: Sample a block $\Theta_{i}$ with probability $p^{+}_{i}$. For each $k\in\Theta_{i}$, set the $k$th entry in an empty row to $\frac{1}{\sqrt{p^{+}_{i}}}$.
@@
Fact A.3 (Matrix Chernoff).
@@
$\operatorname*{\mathbb{E}}[\mathbf{X}_{j}]=\mathbf{I}-\sum_{i=1}^{|\Theta|}p^{+}_{i}\frac{1}{p^{+}_{i}}\sum_{k\in\Theta_{i}}\mathbf{U}_{k}^{\top}\mathbf{U}_{k}=\mathbf{0}$ (19)
@@
$\operatorname*{\mathbb{E}}[\mathbf{X}_{j}^{2}]$ $=\mathbf{I}-2\sum_{i=1}^{|\Theta|}p^{+}_{i}\frac{\sum_{k\in\Theta_{i}}\mathbf{U}_{k}^{\top}\mathbf{U}_{k}}{p^{+}_{i}}+\sum_{i=1}^{|\Theta|}p^{+}_{i}\frac{\left(\mathbf{U}_{\Theta_{i}}^{\top}\mathbf{U}_{\Theta_{i}}\right)^{2}}{{p^{+}_{i}}^{2}}$ (21)
@@
(22$)\preceq n\sum_{i=1}^{|\Theta_{i}|}\frac{2\ell^{\max}_{i}|\Theta_{i}|^{2}\mathbf{U}_{\Theta_{i}}^{\top}\mathbf{U}_{\Theta_{i}}}{\sum_{k\in\Theta_{i}}\ell_{k}}\preceq n\max_{i}2|\Theta_{i}|\frac{\ell^{\max}_{i}}{\ell^{\min}_{i}}\mathbf{I}.$ (23)
Since $|\Theta_{i}|\leq 2$ and the leverage scores in a block are all equal, $\|\operatorname*{\mathbb{E}}[\mathbf{X}_{j}^{2}]\|_{2}\leq 4n$.
Applying Fact A.3 with $m=O(n\log(n/\delta)/\epsilon^{2})$ yields
@@
$\Pr\left(\left\|\mathbf{U^{\top}S^{\top}SV}-\mathbf{U^{\top}V}\right\|_{F}>\epsilon\|\mathbf{U}\|_{F}\|\mathbf{V}\|_{F}\right)$ $\leq\frac{\operatorname*{\mathbb{E}}\left[\left\|\mathbf{U^{\top}S^{\top}SV}-\mathbf{U^{\top}V}\right\|_{F}^{2}\right]}{\epsilon^{2}\|\mathbf{U}\|_{F}^{2}\|\mathbf{V}\|_{F}^{2}}\leq\frac{1}{m\epsilon^{2}}\leq\delta$ (28)
@@
With Lemmas A.2 and A.4 already proved for the special paired leverage score sampling, the following analysis is standard. We include the following proof for completeness.
Proof of Theorem A.1.
@@
$\|\mathbf{A}\tilde{\boldsymbol{\phi}}-\mathbf{b}\|_{2}^{2}=\|\mathbf{A}\tilde{\boldsymbol{\phi}}-\mathbf{A}\boldsymbol{\phi}+\mathbf{A}\boldsymbol{\phi}-\mathbf{b}\|_{2}^{2}=\|\mathbf{A}\tilde{\boldsymbol{\phi}}-\mathbf{A}\boldsymbol{\phi}\|_{2}^{2}+\|\mathbf{A}\boldsymbol{\phi}-\mathbf{b}\|_{2}^{2}$ (29)
where the second equality follows because $\mathbf{A}\boldsymbol{\phi}-\mathbf{b}$ is orthogonal to any vector in the span of $\mathbf{A}$. So to prove the theorem, it suffices to show that
$\|\mathbf{A\tilde{\boldsymbol{\phi}}-A\boldsymbol{\phi}}\|_{2}^{2}\leq\epsilon\|\mathbf{A\boldsymbol{\phi}-b}\|_{2}^{2}.$ (30)
Let $\mathbf{U}\in\mathbb{R}^{\rho\times n}$ be an orthonormal matrix that spans the columns of $\mathbf{A}$. There is some $\mathbf{y}$ such that $\mathbf{Uy}=\mathbf{A}\boldsymbol{\phi}$ and some $\tilde{\mathbf{y}}$ such that $\mathbf{U}\tilde{\mathbf{y}}=\mathbf{A}\tilde{\boldsymbol{\phi}}$. Observe that $\|\mathbf{A\tilde{\boldsymbol{\phi}}-A\boldsymbol{\phi}}\|_{2}=\|\mathbf{U}\tilde{\mathbf{y}}-\mathbf{Uy}\|_{2}=\|\tilde{\mathbf{y}}-\mathbf{y}\|_{2}$ where the last equality follows because $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}$.